OpenStudy (anonymous):

differentiate. f(x)=ln[e^(7x^5)*sqrt(x)]

OpenStudy (anonymous):

im stuck on the first term. can you bring the whoule thing to the front?

OpenStudy (anonymous):

after seperating with the laws and stuff

OpenStudy (anonymous):

is this \[\ln(e^{7x^5}\sqrt{x})\] because if so it is just \[7x^5+\frac{1}{2}\ln(x)\]

OpenStudy (anonymous):

\[\ln(e^{\text{stuff}})=\text{stuff}\]

OpenStudy (anonymous):

oh cuz e and ln cancel? nice.

OpenStudy (anonymous):

i am not sure i would use the work "cancel" but they are inverse functions, so i guess you can think of it that way

OpenStudy (anonymous):

*word